OpenStudy (yisselluna):

Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y? -29/20 31/20 -3/2

OpenStudy (awesomedude2078):

they both have to be positive, because a positive number subtracted from a negative number makes the difference negative

OpenStudy (legomyego180):

you know x = y-3 from the second equation. plug that into the first and solve for y 10(y-3)+10y=1

OpenStudy (3mar):

\[10x+10y=1\] so\[x+y=0.1\] but \[x=y-3\] substitute in the above one \[(y-3)+y=0.1\] \[2y+3=0.1\] \[2y=0.1-3=-2.9\] \[y=-2.9/2=-1.45\] \[y=-\frac{ 29 }{ 20 }\]